OpenStudy (anonymous):

What is the slope of the two points (1, 5) and (6, 4)?

OpenStudy (mimi_x3):

Slope is the gradient so you use the gradient formula \[ m =\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (mimi_x3):

\[(1,5) (6,4) => (x_1,y_1)(x_2,y_2)\]

OpenStudy (mimi_x3):

Are you able to do it now?

OpenStudy (anonymous):

the slope is -1/5

OpenStudy (anonymous):

right

OpenStudy (anonymous):

use the slope formula \[ m \frac{Y_2-Y_1}{X_2-X_1}\] Plugin your points \[ m \frac{4-5}{6-1}\] \[ m -\frac{1}{5}\]

OpenStudy (anonymous):

i got it right :D

OpenStudy (mimi_x3):

yes you are right JackJones.